OpenStudy (anonymous):

how will you determine the equation of the line if the given are two points of the line

OpenStudy (anonymous):

Given points (x1,y1) (x2,y2) Slope=M=(y2-y1)/(x2-x1) Use point-slope form y-y1=m(x-x1) and use algebra to get y=mx+b

OpenStudy (anonymous):

thats actually slope-intercept form but ill but it

OpenStudy (anonymous):

buy

OpenStudy (anonymous):

y-y1=m(x-x1) is point-slope, but after numbers are substituted, it can be rearranged to be in slope-intercept form

Directrix (directrix):

We went through the step by step procedure at this link if anybody wants to view and check it. http://openstudy.com/users/directrix#/updates/4f3a299ee4b0fc0c1a0e7ed7